### v3

# ]https://zf-wei.github.io

Abstract
This study investigates the robustness of graph embedding methods for community detection in the face of network perturbations, specifically edge deletions. Graph embedding techniques, which represent nodes as low-dimensional vectors, are widely used for various graph machine learning tasks due to their ability to capture structural properties of networks effectively. However, the impact of perturbations on the performance of these methods remains relatively understudied. The research considers state-of-the-art graph embedding methods from two families: matrix factorization (e.g., LE, LLE, HOPE, M-NMF) and random walk-based (e.g., DeepWalk, LINE, node2vec). Through experiments conducted on both synthetic and real-world networks, the study reveals varying degrees of robustness within each family of graph embedding methods. The robustness is found to be influenced by factors such as network size, initial community partition strength, and the type of perturbation. Notably, node2vec and LLE consistently demonstrate higher robustness for community detection across different scenarios, including networks with degree and community size heterogeneity. These findings highlight the importance of selecting an appropriate graph embedding method based on the specific characteristics of the network and the task at hand, particularly in scenarios where robustness to perturbations is crucial.
I Introduction
The use of low-dimensional vector representations, or graph embeddings, has garnered significant attention across various disciplines [1, 2, 3]. These representations play a crucial role in characterizing important functional network properties such as robustness and navigability [4, 5, 6, 7]. Additionally, they are instrumental in performing graph analysis tasks including node classification [8], link prediction [9], and visualization [10].
Communities, which are sets of nodes more likely to be connected to each other than to the rest of the graph, are fundamental features of many complex systems represented as graphs [11, 12, 13]. Recent research has demonstrated the potential of using graph embeddings for community detection in networks [14]. This approach leverages the inherent structural information captured in the vector representations to identify cohesive groups of nodes within the network. Overall, graph embeddings offer a versatile framework for analyzing and understanding complex networks, enabling a wide range of applications across various domains.
Community detection using graph embeddings relies on representing graph nodes as points in a low-dimensional vector space, where communities manifest as clusters of points that are close to each other and sufficiently separated from other clusters [14, 15, 16]. These clusters are then recovered using standard data clustering techniques such as $k$-means clustering [17]. Given the practical importance of consistently detecting communities and the effectiveness of graph embedding approaches in doing so, it is essential to investigate the robustness of graph embeddings for community detection under network perturbations. In this context, robustness refers to the ability of a graph embedding method to tolerate perturbations while still accurately recovering the original community structure [18, 19, 20]. This investigation is crucial for understanding how communities identified by graph embedding methods withstand random errors or adversarial attacks, serving as a proxy for evaluating network functionality. Robust graph embeddings ensure the reliability and stability of community detection algorithms in real-world network applications. Moreover, robustness in graph embeddings has practical implications beyond community detection. Robustness to edge deletion in community detection reflects a method’s ability to recover meaningful structure despite incomplete or perturbed network data. In recommendation systems, such robustness ensures that user or item communities remain stable even in the presence of missing interactions, thereby supporting consistent and reliable recommendations. In anomaly detection, it enhances the capacity to distinguish between genuine structural anomalies and incidental perturbations in network communication systems, improving anomaly detection accuracy.
Previous research has investigated the impact of network perturbations on the robustness of community structures using traditional community detection methods [21, 20]. However, little attention has been given to understanding the robustness of graph embedding methods used for community detection under perturbations in the graph topology. While there are numerous graph embedding techniques available for community detection [1, 22], it remains challenging for practitioners to select a robust method and comprehend its implications across various factors, including the size of the network, strength of the original community structure, type of perturbation, and the specific embedding technique used. Moreover, the majority of research on using graph embeddings for community detection has been conducted on synthetic networks, warranting validation of the effects of perturbations on real-world network data. Addressing these gaps in research, we aim to evaluate the robustness of graph embedding methods for community detection under various types of graph topology perturbations, and to identify the methods that demonstrate the highest robustness. By exploring these factors across synthetic and real-world network datasets, we seek to provide insights into the effectiveness and limitations of graph embedding techniques in preserving community structures amidst network perturbations.
In this paper, we address this gap of knowledge by devising a systematic framework to investigate the robustness of graph embedding methods used to extract community structures under network perturbations. We consider both synthetic Lancichinetti-Fortunato-Radicchi (LFR) benchmark graphs [23] and empirical networks with degree and community size heterogeneity.
Our perturbation approach involves the removal of edges based on random node selection [21] and targeted node selection using betweeness centrality [24, 25, 26]. We use seven state-of-the-art graph embedding methods, namely Laplacian eigenmap (LE) [27], locally linear embedding (LLE) [28], higher-order preserving embedding (HOPE) [29], modularized non-negative matrix factorization (M-NMF) [30] in the matrix factorization family; and DeepWalk [31], large-scale information network (LINE) [32], and node2vec [33] in the random walk family. We use $k$-means based on spherical distance [34] to achieve more effective clustering results in high-dimensional spaces. We use element-centric similarity (ECS) [35], a widely adopted community similarity metric, to quantify changes in the community structure after perturbations as a proxy of characterizing graph embedding robustness. This comprehensive methodology allows us to systematically assess and compare the impact of various perturbation strategies on the robustness of graph embedding methods for community detection in both synthetic and real-world networks.
Our findings indicate that incremental selection of nodes and subsequent removal of their adjacent edges reduce ECS. However, this impact tends to be more noticeable across graph embedding methods, but node2vec and LLE, and varies across network sizes, the initial community structure strength, and the perturbation type. This aligns with the expectation that removing a greater number of edges results in a more substantial disruption of the community structure. Notably, targeted node selection leads to a more pronounced and accelerated decline in ECS compared to random node selection. This is rationalized by the fact that adversarial attacks tend to swiftly dismantle the network’s community structure. Among the seven network embedding methods considered, node2vec and LLE consistently prove to be more robust for community detection within their respective graph embedding families; these two methods provide the highest ECS similarity scores. Overall, node2vec is the most robust graph embedding method for community detection. Our results generally concur with those of Kojaku et al. [16], which found that node2vec learns more consistent community structures with networks of heterogeneous degree distributions and size.
We notice that community detection methods based on graph neural networks (GNNs) currently constitute a popular approach [36]. The focus of our study, however, is on shallow graph embedding methods rather than GNNs.
We have made available the code and data to reproduce all the results at  [^1] https://t.ly/uIclh (accessed on August 1, 2025),  [^2] https://t.ly/OzFzV (accessed on August 1, 2025).
II Related Work
II.1 Prior Work Closely Related to the Present Study
There has been substantial research in network community detection leveraging graph embedding techniques.
Tandon et al. [14] tested the ability of the aforementioned seven state-of-the-art graph embedding techniques to detect communities on benchmark graphs. It was reported that if the parameters of the embedding techniques are suitably chosen, the performance of community detection methods based on graph embedding techniques is comparable to the traditional community detection algorithms, including Infomap and Louvain.
Zhang et al. [39] proposed a method that simultaneously performs network embedding and community detection in directed networks. Their model introduced separate vector representations for out-nodes and in-nodes, allowing nodes to belong to different out- and in-communities. Their method has improved performance in community detection while jointly estimating node embeddings and their community assignments. An efficient alternating optimization scheme was developed, and theoretical guarantees for both embedding and community detection are provided, supported by experiments on simulated and real-world networks.
Ghanbari et al. [40] investigated the efficiency of different node embedding algorithms combined with clustering methods for community detection. They evaluated four embedding techniques, including leveraging convolution, attention, inductivity, and shallowness, paired with three clustering algorithms, including KMeans++, DBSCAN, and hierarchical clustering. Experimental results showed that GraphSAGE, an inductive embedding method, combined with KMeans++ produced the best performance.
Zhu et al. [41] proposed a structural equivalence embedding method based on non-negative matrix factorization (SENMF) for community detection, which embeds two similarity measures one capturing node closeness and the other reflecting similarity between distant nodes into a low-dimensional space. Experimental results show that SENMF outperforms several well-known network embedding and traditional community detection methods. In fact, as mentioned in [41], when taking parameters $\theta=0$ and $\varphi=0$, SENMF will equal M-NMF.
Kovács et al. [42] proposed an iterative embedding and reweighting procedure that strengthens intra-community links while weakening inter-community links, enhancing cluster visibility and detectability. This process can lead to such strong geometric separation that even simple link parsing recovers communities with high precision. Additionally, it can serve as a pre-processing step to improve traditional community detection algorithms, including Louvain, Infomap, and HDBSCAN.
Pankratz et al. [43] proposed a method to enhance community detection by initializing heuristic algorithms with partitions obtained from network embeddings. Their approach embeds nodes into a latent space to filter noise while preserving structural relationships, followed by clustering to generate a stable partition. This initialization improves the performance and reduces the volatility of traditional methods such as Louvain and Leiden, particularly in networks with weak community structures. Experimental results demonstrated that the effectiveness of this procedure depends on network topology rather than size.
Zhu et al. [44] proposed a novel Graph Convolutional Network (GCN) method for unsupervised community detection, termed UCDMI. This method leverages mutual information maximization to capture fine-grained information about the global network structure without supervision. A new aggregation function for GCN is introduced to differentiate the importance of neighboring nodes, enhancing the quality of node representations. The proposed approach demonstrates significant improvements in community detection performance through extensive experiments, outperforming several state-of-the-art methods. This method offers a promising direction for unsupervised community detection by integrating mutual information maximization with GCNs.
In this work, we focus on community detection in undirected networks using the seven graph embedding techniques from [14]. Rather than introducing new community detection methods, we examine the robustness of these embedding techniques under network perturbations.
II.2 Other Prior Work Related to the Present Study
The following articles examines the robustness of community structures using traditional community detection methods under different types of network perturbations.
Karrer et al. [18] proposed a framework for evaluating the significance of community structures by perturbing networks with configuration model random graphs and measuring the resulting changes using variation of information. They employed a spectral optimization method for modularity to perform community detection. This approach was applied to a wide range of networks, including both real-world and synthetic examples.
Ramirez-Marquez et al. [45] analyzed the effects of weight uncertainties on community structures in networks with weighted edges. Using Monte Carlo-based uncertainty propagation, they analyzed how variations in link weights influence community structures. The approach was demonstrated through the analysis of an electric power system network, highlighting the significance of weight uncertainties in determining reliable community partitions.
Peche et al. [46] investigated perturbations where the connections between network nodes are influenced by a random geometric graph. They demonstrated that community detection using spectral methods remains robust to such noise, based on a thorough and detailed analysis of the spectrum of the random graph’s adjacency matrix.
Jia et al. [47] proposed a robustness guarantee for community detection against adversarial structural perturbations. For any given community detection method, they introduced a smoothed community detection approach by randomly perturbing the graph structure. They theoretically demonstrated that the smoothed community detection method provably groups a given arbitrary set of nodes into the same community (or different communities) provided that the number of edges added or removed by an attacker is bounded. The method was also empirically evaluated on multiple real-world graphs with ground truth communities.
Tian et al. [20] studied the robustness of community detection algorithms under edge addition perturbations in synthetic and empirical temporal networks. They considered two edge addition scenarios—random and targeted—and evaluated the robustness of communities detected by Infomap, Label Propagation, Leiden, and Louvain algorithms. It was found that robustness performance varies across different community similarity metrics and detection methods, highlighting the dependence of community stability on the choice of algorithm.
Furthermore, given that graph neural network-based community detection methods are numerous and currently constitute a popular approach, to reflect the current research trends, we include the following works on graph community detection using graph neural networks.
Zhang et al. [48] proposed a deep learning-based approach for dynamic community detection that incorporates node relevancy to capture structural information. Additionally, it employs cross-entropy to smooth learned features between adjacent time steps, addressing potential large discrepancies that don’t align with real-world dynamics.
Sun et al. [49] introduced a novel graph neural network encoding method for multiobjective evolutionary algorithms to address community detection in complex attribute networks. The approach transforms the problem into one with continuous decision variables, facilitating smoother fitness landscapes, and demonstrates superior performance on both single- and multi-attribute networks.
[…]
Su et al. [36] provided a comprehensive survey of recent advancements in deep learning techniques for community detection, particularly highlighting their utility in handling high-dimensional network data. It proposes a new taxonomy that categorizes state-of-the-art methods, including various deep neural network models, and summarizes benchmark datasets, evaluation metrics, and open-source implementations.
To address the limitations of previous research, such as the lack of a systematic evaluation of edge removal’s impact on community structure robustness, our study examines edge deletion based on both random and targeted node selection. We apply this approach to both synthetic and real-world networks, accounting for degree and community size heterogeneity. This perturbation approach presents a novel method of network perturbation with real-world relevance. To compare the community structures before and after the network perturbation, we utilized ECS, a robust metric for evaluating network clusters.
III Methods
III.1 LFR Benchmark
Lancichinetti–Fortunato–Radicchi (LFR) benchmark graphs [23, 51] have become a widely adopted tool for generating graphs with ground-truth communities. One of the primary advantages of utilizing the LFR benchmark is its ability to produce graphs where both degree and community size distributions follow power-law distributions. These distributions closely resemble the properties observed in many real-world networks [23, 52, 53, 54]. The exponents governing the degree and community size distributions are controlled by the parameters $\alpha$ and $\beta$, respectively. To generate LFR benchmark graphs, one must specify various other parameters, including the average degree $\langle k\rangle$, maximum degree $maxk$, maximum community size $maxc$, and the mixing parameter $\mu$. The mixing parameter $\mu$ represents the fraction of nodes that share edges across different communities, with lower values of $\mu$ leading to a higher ratio between internal and external edges. This results in increased modularity, a commonly used quality metric for assessing community strength in community research [11]. Thus, smaller values of $\mu$ correspond to stronger partitions. Note that LFR benchmark graphs are undirected and unweighted.
We generate LFR benchmark graphs using the LFR_benchmark_graph method in the publicly available Python package NetworkX [55]. Our experiments involve two different network sizes: 1,0001,0001 , 000 nodes and 10,00010,00010 , 000 nodes, allowing us to investigate the impact of perturbations at varying scales. Furthermore, we explore the influence of the initial community strength by adjusting the mixing parameter, $\mu$. The specific parameter values utilized for the generation of LFR benchmark graphs are detailed in Table 1. Note that the proposed experimental design can be applied to any other synthetic graph generators such as the Girvan–Newman benchmark [56] and the stochastic block model [57].
Table 1: Synthetic experiment parameters.
Parameter Description Value
$N$ number of nodes 1,0001,0001 , 000 and 10,00010,00010 , 000
[…]
$\langle k\rangle$ average degree for LFR 252525
[…]
$\alpha$ degree distribution exponent for LFR 2-2- 2
$\beta$ community size distribution exponent for LFR 1.1-1.1- 1.1
$\mu$ mixing parameter for LFR 0.01,0.1,0.2,0.3,0.4,0.50.01,0.1,0.2,0.3,0.4,0.50.01 , 0.1 , 0.2 , 0.3 , 0.4 , 0.5
$r$ number of realizations 505050
III.2 Edge Removal Procedure
Our edge removal procedure investigates the impact of eliminating edges in the network structure. We employ two distinct methods: edge removal based on random node selection and edge removal based on targeted node selection.
Let’s first discuss edge removal based on random node selection. This simulates random network errors in real-world scenarios [58, 59]. We do so by selecting $p\%$ of nodes randomly and deleting their adjacent edges. To make our experiment unbiased, we independently generate 505050 distinct sets of nodes. In other words, the number of realizations is $r=50$ (see Table 1). We ensure that the removal of edges from each set of nodes does not disconnect the remaining part of the network, which is required by the graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit this set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of the remaining network calculated using data clustering in the embedding space (i.e., $k$-means; see more details in Sec. III.3) and the community structure of the original LFR network using community similarity metrics (i.e., ECS [35]) as detailed in Sec. III.4. The overall similarity score for a given $p\%$ value is calculated by averaging the similarity scores across all 505050 sets.
Second, we discuss targeted edge removal. Targeted edge removal is based on the betweenness centrality of nodes. Betweenness centrality measures the importance of a node in a network based on how often it lies on the shortest paths between other nodes [60]. Nodes with high betweenness centrality usually indicate nodes of vital importance to the network, so removing edges adjacent to these nodes might cripple the functionality of the network [61]. This approach models an adversarial attack in real-world situations [62, 63]. To implement edge removal based on targeted node selection, we first rank the nodes in the network by their betweenness centrality. The node with the lowest betweenness centrality has rank 111, and the node with the highest betweenness centrality has rank $N$, where $N$ is the number of nodes in the network. We then independently generate 505050 sets of nodes, each containing $p\%$ of nodes in the network. The nodes in each set are selected by randomly sampling nodes with probability proportional to their betweenness centrality rank. We ensure that the removal of edges adjacent to each set of nodes does not disconnect the remaining part of the network, which is required by graph embedding methods. In the event that the remaining part of the network becomes disconnected after the removal of a set of edges, we omit the set of nodes and proceed with re-sampling. After each removal, we measure the similarity between the community structure of remaining network computed through data clustering in the embedding space (i.e., $k$-means; see more details in Sec. III.3) and the ground-truth community structure of the original LFR network using ECS. The overall similarity score for a given $p\%$ is calculated by averaging community similarity scores across all 505050 sets.
[…]
III.3 Data Clustering
Community detection (or clustering) is the task of grouping nodes in a network based on their connectivity. A general consensus is that nodes within the same community are more densely connected than nodes across different communities. There are many different community detection algorithms, each with its own strengths and weaknesses [14]. The best algorithm for a particular network depends on the properties of the network and performance constraints [64].
In this article, we perform community detection based on a two-step process consisting of graph embedding [22] and the $k$-means algorithm [65] for data clustering. That is, we control the data clustering step by using the $k$-means method to optimally find the clusters in the low-dimensional representation. Specifically, we perform graph embedding to assign a point in the Euclidean space to each node in the network. Here, we focus on embedding each node of the graph rather than the whole graph (more details on node-embedding methods can be found in [66]). We use seven state-of-the-art graph embedding methods studied in previous research [14] across two different families of graph embedding methods, including: (1) matrix-based methods and (2) random walk methods (according to the taxonomy by M. Xu in [22]). In the matrix-based family, we include Laplacian eigenmap (LE) [27], locally linear embedding (LLE) [28], higher-order preserving embedding (HOPE) [29], and modularized non-negative matrix factorization (M-NMF) [19]. In the random walk family, we include DeepWalk [31], large-scale information network (LINE) [32], and node2vec [33]. Note that we make the distinction between graph embedding families to stress the foundational mechanism under which graph embeddings are constructed. We, however, note that under specific conditions regarding the window size of the skip-gram model, random walk methods are equivalent to matrix factorization methods [67]. Tandon et. al provides a thorough test on graph embedding methods for data clustering, which have been shown to perform well for community detection in networks [14].
We use the publicly available Python package nxt_gem for LE and HOPE [68]. As there is no correct implementation of LLE for graphs available online, we developed our own implementation available on GitHub in  [^3] https://github.com/zf-wei/LLECupy (accessed on April 29, 2024). The package karateclub includes implementations of DeepWalk and M-NMF [70]. The package GraphEmbedding for LINE is publicly available on GitHub in  [^4] https://github.com/shenweichen/GraphEmbedding (accessed on April 29, 2024). We utilize node2vec package for node2vec embedding available online  [^5] https://pypi.org/project/node2vec (accessed on April 29, 2024).
After computing graph embeddings, we use the $k$-means algorithm to group the points in the Euclidean space into clusters. Note that other data clustering algorithms used before to cluster embedding results, such as Gaussian mixture models [73], lead to similar results as $k$-means [14]. The $k$-means algorithm, a popular method in data clustering, minimizes the distance between each data point and its centroid. It works by iteratively assigning data points to the nearest cluster centroid and then updating the centroids based on the average of the assigned points until convergence, aiming to minimize the within-cluster variance. In our case, the data points are the embedding points in Euclidean space for each network node and centroids are virtual points representing the clusters. As in high-dimensional spaces, the concept of proximity may not be meaningful as different distances may select different neighbors for the same point [74]; we use $k$-means based on spherical distance to get better clustering outcome. According to [34], $k$-means based on spherical distance is a widely used clustering algorithm for sparse and high-dimensional data. Specifically, for two nodes embedded as points $\mathbf{x}=(x_{1},x_{2},\ldots,x_{d})$ and $\mathbf{y}=(y_{1},y_{2},\ldots,y_{d})$ in the $d$-dimensional Euclidean space, the spherical distance between the two nodes is defined as
$\mathsf{d}_{S}(\mathbf{x},\mathbf{y})=1-\frac{\sum_{i=1}^{d}x_{i}y_{i}}{\sqrt{\sum_{i=1}^{d}x_{i}^{2}\sum_{i=1}^{d}y_{i}^{2}}}.$
As data clustering methods often necessitate prior knowledge of the number of clusters, we provide the data clustering procedure with the correct number of clusters derived from the LFR benchmark graphs. By doing so, we ensure a fair evaluation of the embedding methods. Note that such a luxury may be absent when analyzing real-world networks, as the true number of clusters is typically unknown in those cases.
III.4 Community Similarity
To measure the similarity between community partitions, we employ ECS [35], computed with the CluSim package [75]. Compared to other clustering similarity metrics such as the normalized mutual information (NMI) [76], ECS is a most robust metric for comparing clusters and better addresses challenges such as biases in randomized membership, skewed cluster sizes, and the problem of matching [35]. We use default parameter values in ECS computation. ECS yields values in the range from 0 to 111, where a higher value indicates a greater similarity between partitions.
Fig. 1 is a schematic diagram showing the process we follow in our experimental design.
Network Generation, Sec. III.1Perturbation, Sec. III.2Data Clustering, Sec. III.3Similarity, Sec. III.4
Figure 1: Schematic diagram of experimental design.
IV Results and Discussion
In this section, we present the results of our experiments. Recall that Sec. III.3 details the embedding methods we use here. In the following subsections, we use $\mathbf{A}$ to denote the adjacency matrix of a graph. We report results based on the family of graph embedding methods (i.e., matrix-based family in Sec. IV.1 and random walk family in Sec. IV.2).
In the subsequent line figures, unless otherwise stated (in Figs. 17, 18 and 14), subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on small networks with 1,0001,0001 , 000 nodes and subfigure (b) depicts experimental results on large networks with 10,00010,00010 , 000 nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection, for small and large networks respectively.
We present line figures using 161616-dimensional and 323232-dimensional embedding. We focus only in these two dimensions as smaller embedding dimensions seem to be enough to provide an accurate low-dimensional vector representation of networks [15, 16]. Recall that a higher ECS value indicates a greater similarity between network community partitions obtained from the LFR benchmark graphs and the embeddings. Thus, when analyzing line figures, higher ECS values and therefore higher positions of line curves indicate a more robust graph embedding method for community detection.
Table 2 shows the experimental parameters employed by the embedding methods, excluding embedding dimension, $d$. Notably, LE and LLE methods only require the embedding dimension $d$. For the remaining parameters, we utilized the default parameters as specified in Tandon et al. (2021) [14]; they show that default values of parameters for the embedding methods lead to comparable performance as traditional community detection methods, and that the optimized parameters do not exhibit a significant improvement over default parameters.
Table 2: Parameter values for embedding methods.
Family Method Parameters
Matrix HOPE $\beta=0.01$
Matrix M-NMF $\lambda=0.2$, $\alpha=\beta=0.05$, $\eta=5$, $N=200$
Random Walk DeepWalk $t=40$, $w=10$, $n=80$
Random Walk LINE batch size: 819281928192, epochs: 100100100
Random Walk node2vec $t=10$, $w=10$, $n=80$, $p=q=1$
Finally, Sec. IV.3 presents experimental results to evaluate the robustness of our community detection methods on two real-world networks with labeled communities: the email-EU-core network [77] and the AS network [7]. These networks exhibit similar characteristics to synthetic counterparts, specifically power-law degree and community size distributions as observed in LFR benchmark graphs. Our experimental methodology and insights for real-world networks mirror findings for synthetic counterparts.
IV.1 Matrix Factorization Methods
Matrix-based methods project network nodes into a Euclidean space via eigenvectors following a similar idea as in spectral clustering [78]. In this family of methods, we include Laplacian eigenmap (LE) [1, 27], locally linear embedding (LLE) [14, 28], higher-order preserving embedding (HOPE) [1, 29], and modularized non-negative matrix factorization (M-NMF) [14, 19]. We now detail results in each of them.
IV.1.1 Laplacian Eigenmap (LE)
Laplacian eigenmap (LE) [1, 27] aims to minimize the objective function
$\sum_{i,j}\left|\mathbf{x}_{i}-\mathbf{x}_{j}\right|^{2}\mathbf{A}_{ij},$
subject to the constraint $\mathbf{x}^{T}\mathbf{D}\mathbf{x}=\mathbf{I}$, where $\mathbf{D}$ represents the diagonal matrix of graph node degrees, and $\mathbf{x}_{i}$ denotes the vector indicating the position of the point representing node $i$ in the embedding. For $d$-dimensional embedding with LE, the solution can be obtained by extracting the eigenvectors corresponding to the $d$ smallest eigenvalues (except the zero eigenvalue) of the normalized Laplacian matrix
$\mathbf{D}^{-1}(\mathbf{D}-\mathbf{A}).$
Our experimental results on the impact of edge removal, generated through the LE embedding method with dimension 161616, are reported in Fig. 2. Our experimental results using LE embedding method with dimension 323232 are collected in Fig. 3.
Figure 2: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with 1,0001,0001 , 000 nodes, and (b) and (d) are for networks with 10,00010,00010 , 000 nodes.
Figure 3: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with 1,0001,0001 , 000 nodes, and (b) and (d) are for networks with 10,00010,00010 , 000 nodes.
[…]
IV.1.2 Locally Linear Embedding (LLE)
Locally linear embedding (LLE) [14, 28] minimizes the objective function
$\sum_{i}\Bigl{|}\mathbf{x}_{i}-\sum_{j}\mathbf{A}_{ij}\mathbf{x}_{j}\Bigr{|}^{2}.$
Given the form of the objective function, each point in the embedded space $\mathbf{x}_{i}$ is approximated as a linear combination of its neighbors in the original graph. To ensure a well-posed problem, solutions are required to be centered at the origin, i.e., $\sum_{i}\mathbf{x}_{i}=0$, and to have unit variance, i.e., $\frac{1}{N}\mathbf{x}^{T}\mathbf{x}=\mathbf{I}$. Subject to these constraints, for a $d$-dimensional embedding with LLE, the solution is approximated by the eigenvectors corresponding to the lowest $d$ eigenvalues (excluding the zero eigenvalue) of the matrix $(\mathbf{I}-\mathbf{A})^{T}(\mathbf{I}-\mathbf{A})$.
Our experimental results on the impact of edge removal, generated through the LLE embedding method with dimensions 161616 and 323232, are reported in Fig. 4 and Fig. 5, respectively.
Figure 4: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have 1,0001,0001 , 000 nodes, and those in (b) and (d) have 10,00010,00010 , 000 nodes.
Figure 5: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have 1,0001,0001 , 000 nodes, and those in (b) and (d) have 10,00010,00010 , 000 nodes.
The insights drawn from Figs. 4 and 5 for LLE align closely with the insights derived for LE. In fact, the curves in Fig. 4 are slightly higher in position than those in Fig. 2, indicating that LLE outperforms LE. For example, the data points on the green curve (corresponding to $\mu=0.4$) from LLE in Fig. 4(a) have higher ECS scores over LE in Fig. 2(a) for a given percentage of selected nodes.
IV.1.3 Higher-Order Preserving Embedding (HOPE)
Higher-order preserving embedding (HOPE) is designed to maintain the similarity between nodes [1, 29]. The objective function being minimized is $\bigl{|}\mathbf{S}-\mathbf{x}\mathbf{x}^{T}\bigr{|}$, where $\mathbf{S}$ is the Katz similarity matrix defined as
$\mathbf{S}=\beta\sum_{\ell=1}^{\infty}\mathbf{A}^{\ell},$
with $\beta$ denoting the decay parameter. In our experiments, we set $\beta=0.01$.
Our experimental results on the impact of edge removal, generated through the HOPE embedding method with dimensions 161616 and 323232, are reported in Fig. 6 and Fig. 7, respectively.
Figure 6: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain 1,0001,0001 , 000 nodes, while those in (b) and (d) consist of 10,00010,00010 , 000 nodes.
Figure 7: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain 1,0001,0001 , 000 nodes, while those in (b) and (d) consist of 10,00010,00010 , 000 nodes.
For the HOPE embedding method, our results suggest that using a 32-dimensional embedding (i.e., Fig. 7) brings higher ECS scores than a 16-dimensional one (i.e., Fig. 6). For example, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 7(a) have higher ECS scores than those data points on the green curve (corresponding to $\mu=0.4$) in Fig. 6(a) at respective percentage of selected nodes. We can also see that HOPE has higher ECS scores with smaller graphs.
IV.1.4 Modularized Non-Negative Matrix Factorization (M-NMF)
Modularized Non-Negative Matrix Factorization (M-NMF) [19] preserves both the first-order proximity (pairwise node similarity) and the community structure for network embedding. Modularity, which evaluates the quality of community partitions within a network by comparing the observed network with randomized versions that lack inherent community structure, is integrated into the optimization function of M-NMF. While minimizing the optimization function of M-NMF (see [19]), nodes are brought into proximity when they exhibit similarity and simultaneously when they are part of clusters derived from high-modularity partitions [14].
The following parameter values are used in our experiments: $\lambda=0.2$, $\alpha=0.05$, $\beta=0.05$, $\eta=5.0$, and number of iterations $N=200$.
Our experimental results on the impact of edge removal, generated through the M-NMF embedding method with dimensions 161616 and 323232, are reported in Fig. 8 and Fig. 9, respectively.
Figure 8: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
Figure 9: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
From Figs. 8 and 9, we can see that M-NMF produces higher ECS scores with smaller graphs.
Table 3: Choice of more robust graph embedding methods in the family of matrix-factorization.
Weaker Community Structure (larger $\mu$) Stronger Community Structure (smaller $\mu$)
Smaller Network (smaller $N$) LLE with a low embedding dimension LLE with a low embedding dimension, or HOPE with a high embedding dimension
Larger Network (larger $N$) LE with a higher embedding dimension LE or LLE with a lower embedding dimension
IV.1.5 Summary of Matrix Factorization Methods
[…]
For smaller real-world graphs containing a few thousand vertices, to obtain more robust partitions, the LLE method is suggested with a modest embedding dimension, typically on the order of tens, such as 16 dimensions; Alternatively, when the community partition is stronger, the HOPE embedding method with a higher embedding dimension produces comparably robust partitions. In the case of larger real-world graphs with a weaker community partition, to obtain more robust community partitions, it is suggested to use the LE embedding method with a moderately higher embedding dimension, such as 323232 dimensions as utilized in our experiments (see Fig. 3). When dealing with larger real-world graphs characterized by a stronger community structure, the LLE or LE embedding method with a lower embedding dimension produces more robust community partitions. In summary, LLE is the most robust method among matrix factorization methods. Table 3 summarizes our findings for matrix factorization methods.
IV.2 Random Walk Methods
Random walk methods learn network embeddings of graph nodes by modeling a stream of short random walks. We chose DeepWalk [31], LINE [32], and node2vec [33] as widely used methods of this family.
IV.2.1 DeepWalk
DeepWalk [31] extends language modeling techniques to graphs, departing from words and sentences. This algorithm leverages local information acquired through random walks, treating these walks as analogous to sentences in the word2vec [79] language modeling approach. To generate a random walk originating from a specified starting node, neighbors of the current node in the walk are randomly selected and added to the walk iteratively until the intended walk length is achieved.
In our experiments, we use the following parameters: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$.
Our experimental results on the impact of edge removal, generated through the DeepWalk embedding method with dimension 161616, are reported in Fig. 10. Our experimental results using DeepWalk embedding method with dimension 323232 are collected in Fig. 11.
Figure 10: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings.Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with 1,0001,0001 , 000 nodes, and (b) and (d) represent those with 10,00010,00010 , 000 nodes.
Figure 11: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with 1,0001,0001 , 000 nodes, and (b) and (d) represent those with 10,00010,00010 , 000 nodes.
From Figs. 10 and 11 for DeepWalk, we can observe that networks with stronger initial partitions (i.e., lower $\mu$ values) experience less perturbation impact and targeted node selection has a significant influence on clusters. Moreover, from the right column of Fig. 11, we see that some curves exhibit unexpected increase as we progressively selected more nodes.
IV.2.2 Large-Scale Information Network Embedding (LINE)
The large-scale information network embedding (LINE) method aims to position nodes in close proximity to each other as their similarity increases. LINE can be implemented based on first-order similarity, second-order similarity, or both. According to [14], it works the best when implemented based on first-order similarity, i.e., adjacency of nodes. In fact, for unweighted networks, the LINE method based on first-order similarity minimizes this objective function:
$-\sum_{(i,j)\in E}\log\Bigl{(}\frac{1}{1+\exp{(-\mathbf{x}_{i}\cdot\mathbf{x}_{j})}}\Bigr{)},$
where $E$ is the edge set and $\mathbf{x}_{i}$ denotes the embedding point of node $i$.
The optimization process of the LINE method utilizes stochastic gradient descent, which is enhanced by an edge-sampling treatment, as detailed in [32]. Specifically, the utilization of the LINE embedding method necessitates the presence of GPU for computation. In our study, the machine learning model was trained using a batch size of 819281928192 over 100100100 epochs, facilitating iterative updates to the embedding vector across the entire training dataset.
Our experimental results on the impact of edge removal, generated through the LINE embedding method with dimensions 161616 and 323232, are reported in Fig. 12 and Fig. 13, respectively.
Figure 12: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with 1,0001,0001 , 000 nodes correspond to (a) and (c), and those with 10,00010,00010 , 000 nodes correspond to (b) and (d).
Figure 13: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with 1,0001,0001 , 000 nodes correspond to (a) and (c), and those with 10,00010,00010 , 000 nodes correspond to (b) and (d).
From Figs. 12 and 13. the LINE embedding method produces higher ECS scores when applied to larger networks. Our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals that curves exhibit a steep decline followed by a subsequent increase, but we notice that increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase, rendering the pattern less pronounced, as indicated in Fig. 14. Based on the aforementioned observation, LINE becomes a viable option for network clustering when dealing with relatively larger networks and the computational resources, particularly GPUs, are available.
Figure 14: Mean ECS using LINE method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with 1,0001,0001 , 000 nodes. The first plot displays results using 128128128-dimensional embeddings; the second plot displays results using 256256256-dimensional embeddings.
IV.2.3 node2vec
node2vec employs a similar optimization procedure to DeepWalk, but the process of generating “sentences” differs [33]. Specifically, simple random walk is used for DeepWalk, while biased random walk is utilized for node2vec. These biased random walks are composed of a blend of steps following both breadth-first and depth-first search strategies, with parameters $p$ and $q$ controlling the respective influences of these two strategies (see [33] for details about how random walk are generated in node2vec). Therefore, node2vec produces higher-quality and more informative embeddings than DeepWalk.
We use the following parameters in our experiments: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$, biased walk weights $p=1$ and $q=1$.
Our experimental results on the impact of edge removal, generated through the node2vec embedding method with dimensions 161616 and 323232, are reported in Fig. 15 and Fig. 16, respectively.
Figure 15: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with 161616-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with 1,0001,0001 , 000-node networks, and (b) and (d) with 10,00010,00010 , 000-node networks.
Figure 16: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with 1,0001,0001 , 000-node networks, and (b) and (d) with 10,00010,00010 , 000-node networks.
In the case of node2vec, we consistently observe a decrease in ECS as we iteratively select additional nodes and remove their adjacent edges, which aligns with our intuitive expectations.
IV.2.4 Summary of Random Walk Methods
Similar to matrix factorization methods, the curve tends to exhibit higher values for smaller mixing parameters ($\mu$). Moreover, when using targeted node selection, the ECS shows a more pronounced and rapid decline compared to random node selection, as evident from the comparison between Fig. 15(a) and Fig. 15(c).
From our discussion of each method, we observe that DeepWalk may exhibit an unexpected increase while LINE experiences unexpected oscillations. In contrast, node2vec is free from these issues and consistently achieves the highest ECS similarity scores. See the Appendix A for a brief explanation on the advantage of Node2vec embedding in network community detection.
When considering the application of random walk embeddings for community detection, node2vec consistently proves to be a more robust choice over DeepWalk and LINE, across networks of varying sizes, the strength of the initial community partition, and perturbation type.
IV.3 Real-World Networks
In this section, we present the results of our experiments conducted on two real-world networks with labeled communities, i.e., the email-EU-core network and the AS network. The email-EU-core network is constructed using email data sourced from a prominent European research institution, comprising 986986986 nodes and 16,68716,68716 , 687 edges. For this network, we use the departmental affiliations of its members as the underlying community structure [77]. The AS network, derived from the AS Internet topology data collected in June 2009 by the Archipelago active measurement infrastructure, represents the inter-domain Internet topology with Autonomous Systems (ASes) as nodes and AS peerings as links, forming an AS-level topology graph. This dataset comprises 23,74823,74823 , 748 nodes and 58,41458,41458 , 414 edges. For this network, we use the densely connected AS groups as the underlying community structure [7]. We select them because they closely resemble power-law degree and community size distributions, as the synthetic networks counterpart (i.e., LFR benchmark graphs). To confirm the alignment of these real-world networks with the properties of LFR networks, we conducted statistical tests using the Python package powerlaw based on the statistical test derived in the work by Clauset et al [80]. The results indicate that both the email-EU-core network and the AS network exhibit power-law characteristics [81]. Specifically, in the email-EU-core network, the degree distribution adheres to a power-law distribution and the community size sequence follows an exponential distribution; while in the AS network, both degree and community size distributions conform to power laws.
[…]
We report experimental results using different methods with 161616-dimensional embedding on these two real-world networks in Fig. 17. The experimental results with 323232-dimensional embedding on the two real-world networks are depicted in Fig. 18.
Figure 17: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with 161616-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
Figure 18: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with 323232-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
[…]
As the number of removed selected nodes increases, the ECS sometimes rises, for example, in Fig. 17(a) and Fig. 18(a) for the M-NMF mehtod. To get more insights and mitigate this pathological increasing behavior, we have performed additional experiment and analyzed the ECS score of M-NMF with 128-dimensional embeddings. We notice that for clustering bigger networks where the expected number of communities is naturally higher, M-NMF required the use of a higher embedding dimension to effectively separate those communities. See the Appendix B for more details.
For the two real-world networks we analyzed, the “underlying community structure” is assigned based on the functional roles and real-world context of each network. While this assignment is somewhat artificial, we believe it is meaningful and provides a reasonable approximation. By leveraging these predefined community structures, we can still gain valuable insights into the robustness of our community detection methods using network embeddings.
IV.4 Quantitative Comparison of Graph Embedding Methods
[…]
$\Delta y=\frac{1}{13}\sum_{i=1}^{13}|y_{i}-y_{i-1}|.$
This quantity captures the average rate of decline in similarity as more nodes are selected and their adjacent edges removed. A smaller $\Delta y$ indicates that the embedding method better preserves the community structure under perturbation, i.e., it is more robust. Naturally, a more robust embedding method is characterized by having a larger $y_{0}$ (indicating better initial similarity) and a smaller $\Delta y$ (indicating better capacity to “maintain” community structure as more nodes are selected). Table 4 shows the numbers we got for the aforementioned parameters:
Table 4: Mean initial ECS values ($y_{0}$) and average decay rates ($\Delta y$) for different graph embedding methods under edge-deletion perturbations. Higher $y_{0}$ and lower $\Delta y$ indicate more robust community structure preservation. Results are based on networks with 10,00010{,}00010 , 000 nodes, embedding dimension 323232, and mixing parameter $\mu=0.1$.
Family Matrix Factorization Methods Random Walk Methods
Methods LE LLE HOPE M-NMF DeepWalk LINE node2vec
$y_{0}$ 0.914 1.000 0.566 0.943 0.853 0.881 1.000
$\Delta y$ 0.012 0.005 0.027 0.022 0.007 0.012 0.001
As observed from Table 4, for the parameter choices mentioned above, LLE and node2vec provide the most robust community detection results. This verifies our qualitative analysis. Additionally, we explore networks with varying node counts (1,0001,0001 , 000 or 10,00010,00010 , 000) and different mixing parameters ($\mu=0.01,0.1,0.2,0.3,0.4,0.5$), and we also conducted experiment with embedding dimensions of 161616 or 323232, for both random and targeted node selection. Given the variety of parameters, we compute for each embedding method the mean values over these configurations. We summarize the overall robustness of community detection across these seven graph embedding methods in Table 5.
Table 5: Mean robustness metrics for different embedding methods, averaged over multiple experimental settings including varying network sizes (1,0001{,}0001 , 000 and 10,00010{,}00010 , 000 nodes), mixing parameters ($\mu=0.01$ to 0.50.50.5), embedding dimensions (161616 and 323232), and both random and targeted node selection strategies.
Family Matrix Factorization Methods Random Walk Methods
Methods LE LLE HOPE M-NMF DeepWalk LINE node2vec
$y_{0}$ 0.896 0.912 0.651 0.766 0.837 0.877 0.921
$\Delta y$ 0.026 0.025 0.022 0.029 0.019 0.040 0.017
From these tables, we can confirm that LLE and node2vec provide the most robust community detection results, which aligns with our qualitative analysis in the manuscript. Note that, in Table 5, LLE has larger $\Delta y$ than HOPE; however, since LLE has a much larger $y_{0}$ than HOPE, we still identify LLE as the best method within the family of matrix-factorization methods. To make it more visually convenient for the readers, we have visualize Table 5 as Fig. 19. In Fig. 19, points near the bottom-right corner correspond to more robust embedding methods. Again, LLE and node2vec outperform other embedding methods and show best overall robustness.
Figure 19: Visualization of overall robustness metrics ($y_{0}$ and $\Delta y$) for different embedding methods, averaged across all configurations including both random and targeted node selection, various network sizes, mixing parameters, and embedding dimensions.
IV.5 Limitations of Our Study
Having discussed the insights derived from our results, we acknowledge the following limitations of our work:
  1. 1.
    Our experiments focus solely on LFR benchmark graphs, while other benchmarks, such as stochastic block models random graphs, remain unexplored.
  2. 2.
    Our study emphasizes computational investigations of community detection robustness rather than analytical research. Consequently, we do not analyze detectability limits, which remain underexplored for LFR benchmark graphs due to their complex generation and structural properties.
  3. 3.
    In our experiments on two real-world networks, the underlying community structure of ground truth of communities was assigned based on real-world context as determined by domain experts. However, such assignments may introduce subjectivity and potential biases.
  4. 4.
    When we study the two real-world networks, we have artificially reconnecting disconnected components by introducing random edges. This was necessary due to a technical limitation of the graph embedding methods we employed, which tend to perform poorly on disconnected graphs. In fact, the graph embedding methods we used embed each connected component independently for disconnected graphs, leading to arbitrary relative positioning between components, which undermines the global consistency needed for downstream tasks such as clustering or community detection. We acknowledge that artificial reconnection would introduce some deviation from the original network structure.
V Conclusion
We conducted a systematic set of experiments aimed at evaluating the robustness of graph embedding methods for community detection on networks. Our study encompassed LFR benchmark graphs and two real-world networks, with variations in network sizes and mixing parameters considered for the LFR benchmark graphs. We studied two different strategies for perturbations: (1) edge removal based on random node selection, analogous to random network errors, and (2) edge removal based on targeted node selection, modeling deliberate attacks. Both perturbation strategies involve the removal of adjacent edges after selecting nodes. Our community detection approach utilized seven graph embedding methods. Community similarity was measured through ECS.
Our experimental findings revealed a general decline in community similarity as more nodes were selected and their adjacent edges deleted. Notably, targeted node selection led to a more pronounced and rapid decline in community similarity compared to random node selection. Analysis of clustering similarity scores suggested that LFR networks with lower mixing parameters, indicative of a stronger community structure, exhibit greater robustness to network perturbation. Moreover, diverse graph embedding methods displayed varying degrees of robustness in network community detection. Specifically, LLE demonstrated proficiency within the family of matrix-based methods. For smaller real-world networks with a limited number of vertices, the more robust embedding method is LLE with a modest embedding dimension, typically on the order of tens — such as the 16 dimensions employed in our experiment. In the case of larger real-world networks, the LE and LLE embedding methods showed greater robustness to partition. Specifically, LE with a higher embedding dimension is preferable when the community structure is less pronounced, whereas LE or LLE with a lower embedding dimension is more effective when the network exhibits a stronger community structure. Within the family of random walk methods, our results consistently highlighted node2vec as the superior choice. Notably, node2vec also outperformed all methods within the family of matrix factorization methods.
In contemplating potential avenues for future research, it is noteworthy that our current experiments did not involve parameter optimization in the graph embedding algorithms, primarily due to its inherent computational expense. To attain a more exhaustive and comprehensive insight into the robustness of community detection using graph embedding methods, we may explore the incorporation of parameter optimization in our forthcoming studies. Moreover, while this work focuses on shallow graph embeddings, we acknowledge that neural network-based methods have shown significant promise in recent years. GNNs, particularly in clustering tasks, leverage node features and graph topology in an end-to-end learning framework, providing powerful embeddings that can capture complex relationships within graphs. We plan to explore such approaches in future work to compare their robustness with traditional methods.
Acknowledgements.
The work of Zhi-Feng Wei is supported by the Department of Energy (DOE) Office of Advanced Scientific Computing Research (ASCR) through the ASCR Distinguished Computational Mathematics Postdoctoral Fellowship. Pacific Northwest National Laboratory (PNNL) is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830. The research of Zhi-Feng Wei was supported in part by an appointment with the National Science Foundation (NSF) Mathematical Sciences Graduate Internship (MSGI) Program. This program is administered by the Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy (DOE) and NSF. ORISE is managed for DOE by ORAU. All opinions expressed in this paper are the author’s and do not necessarily reflect the policies and views of NSF, ORAU/ORISE, or DOE. The research of Zhi-Feng Wei was supported in part by an appointment to the Oak Ridge National Laboratory GRO Program, sponsored by the U.S. Department of Energy and administered by the Oak Ridge Institute for Science and Education. This paper has been coauthored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The publisher, by accepting the article for publication, acknowledges that the U.S. government retains a nonexclusive, paid up, irrevocable, world-wide license to publish or reproduce the published form of the manuscript, or allow others to do so, for U.S. government purposes. The DOE will provide public access to these results in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.
Appendix A Remarks on node2vec method
Node2vec produces more robust graph community detection results due to its flexible random walk strategy, which employs biased random walks to provide a trade-off between breadth-first search (BFS) and depth-first search (DFS). As mentioned in [1], node2vec produces higher-quality and more informative embeddings than DeepWalk, enabling node2vec to preserve community structure as well as structural equivalence between nodes.
There has been theoretical evidence supporting the superiority of node2vec over DeepWalk in community detection tasks. As demonstrated in [16, 82], node2vec is notably robust to degree heterogeneity due to its ability to learn degree-agnostic embeddings. In contrast, DeepWalk tends to encode node degree as a dominant feature in the embedding space. This sensitivity makes DeepWalk more susceptible to structural noise in networks with heterogeneous degree distributions, such as those generated by the LFR benchmark model. These theoretical insights are consistent with our empirical results. As shown in Fig. 10 and Fig. 15, when the proportion of selected nodes is small—so the perturbed network retains structural properties close to the original LFR benchmark—the ECS similarity scores achieved by DeepWalk are consistently lower than those by node2vec. A similar pattern is observed in Fig. 11 and Fig. 16. In addition, this theoretical insight is reflected in Tables 4 and 5: the $y_{0}$ values for DeepWalk are smaller than those for node2vec in these two tables.
Furthermore, Cohen et al. [83] studied the robustness of networks with power-law degree distributions under random breakdowns using percolation theory. Specifically, they showed that if a network has a degree distribution with probability mass function $\{P_{k}\}$, and a fraction $p$ of its nodes break down, the resulting degree distribution has probability mass function
$P^{\prime}_{k}=\sum_{\ell=k}^{\infty}P_{\ell}\binom{\ell}{k}(1-p)^{k}p^{\ell-k}.$
[…]
$\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[S_{N}>x\right]\sim\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[N>\tfrac{x}{1-p}\right],\quad\text{as }x\to\infty.$
Because $N$ is a power-law random variable with exponent $\gamma$ , it follows that
$\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[S_{N}>x\right]\sim\bigl{(}\tfrac{x}{1-p}\bigr{)}^{-\gamma},\quad\text{as }x\to\infty.$
Now that $\{P^{\prime}_{k}\}$ is the probability mass function of $S_{N}$, we proved our assertion: if $\{P_{k}\}$ is power-law distributed with exponent $\gamma$, then $\{P^{\prime}_{k}\}$ is asymptotically power-law distributed with the same exponent $\gamma$. Hence, the type of random perturbation we considered do not eliminate degree heterogeneity. Consequently, DeepWalk remains vulnerable to degree-based noise after such perturbations, while node2vec continues to perform robustly in community detection.
As discussed in [16, 67], LINE can be regarded as a special case of node2vec with a window size $w=1$, meaning that it focuses solely on learning dyadic (i.e., direct) relationships between nodes. This limited temporal context restricts LINE’s ability to capture higher-order or indirect structural features. Consequently, LINE exhibits lower robustness than node2vec, particularly in noisy or perturbed networks, because it tends to encode stochastic and noisy pairwise interactions. In contrast, node2vec, by aggregating longer biased random walks, can effectively capture indirect relationships between nodes, making it more resilient in downstream tasks such as community detection.
Moreover, as mentioned in [16], node2vec is robust to network sparsity compared to traditional spectral embedding methods. These observations are corroborated in our experiments. In fact, as we select more nodes and delete their adjacent edges, networks are getting sparsified. Therefore, it is understandable that node2vec is more robust in our experiments.
Hence, node2vec is more robust against random and targeted perturbations in community detection, outperforming other methods in stability under perturbations.
Appendix B M-NMF method with higher embedding dimensions
An increase in ECS can be observed in Figs. 9, 17 and 18 for the M-NMF method, where 16-dimensional and 323232-dimensional embedding are used.
We see some unexpected increasing behaviors in the M-NMF curves in Fig. 9(b) and Fig. 9(d) for synthetic LFR networks with 10,00010,00010 , 000 nodes. Intuitively, these networks are bigger and therefore more complex, which makes it more difficult to detect communities. We thus conducted additional experiments on M-NMF with a 128128128-dimensional embedding as shown in Fig. 20. We observe that the unexpected increasing behaviors are mitigated compared to 323232-dimensional embedding. Specifically, the ECS scores of M-NMF with 128128128-dimensional embedding are more stable than those with 323232-dimensional embedding. This indicates that the choice of embedding dimension can significantly and sensitively impact the performance of M-NMF. Due to the substantial computational cost associated with our additional experiments on synthetic networks with 10,00010,00010 , 000 nodes, we reduced the number of realizations per percentage of selected nodes to 101010, compared to the 505050 used in the original paper, to obtain results within the high-performance computing cluster’s time limit. Standard deviations are still negligible compared to the mean but we do not show them here.
Figure 20: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 128128128-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
Notably, the M-NMF curves in Fig. 17(a) and Fig. 18(a) for email-EU-core networks are generated using the exact same node selection sequence, with the only difference being the embedding dimension. In Fig. 17(a), where the embedding dimension is 161616, the M-NMF curve remains almost stable showing slightly unexpected increases. In Fig. 18(a), the increasing of ECS is more noticeable. To get more insights into this behavior, we performed additional experiment and analyzed the ECS scores of M-NMF with 128128128-dimensional embedding. The comparison results are presented in Fig. 21. We can see that the ECS scores of M-NMF with 128128128-dimensional embedding are more “stable” as more nodes are selected, where unexpected increasing behaviors are much less pronounced. This again suggests that the choice of embedding dimension can significantly impact the performance of M-NMF, as it has been noticed by [14] and [41]. Furthermore, we observe a difference in community structure: our synthetic LFR networks with 1,0001,0001 , 000 nodes exhibit approximately 101010 communities; our synthetic LFR networks with 10,00010,00010 , 000 nodes exhibit approximately 202020 communities. In contrast, the email-EU-core network contains 424242. Consequently, employing a higher embedding dimension, such as 128128128, appears reasonable to effectively separate these more numerous communities in the embedding space for email-EU-core network.
Figure 21: Mean ECS using M-NMF method on email-EU-core network with 986986986 nodes with different embedding dimensions.
In summary, we have observed that the choice of embedding dimension can significantly sensitively impact the performance of M-NMF in community detection, which can lead to pathological behaviors in the ECS scores. This reveals that M-NMF is not as robust compared to other methods, especially node2vec, which is more stable and less sensitive to the choice of embedding dimension.